OpenStudy (vegan555):

equation of a parabola: 3x^2-6(x+y)=10 Find Vertex, Focus, and Directrix

OpenStudy (campbell_st):

well it'd rewrite it as \[3x^2 -6x -6y = 10~~~ or ~~~6y = 3x^2 - 6x - 10\] then I'd look at factoring the right hand side \[6y = 3(x^2 - 2x) - 10\] and complete the square in x \[6y = 3(x^2 -2x + 1) - 10 - 1\] which becomes \[6y = 3(x - 1)^2 - 11\] I'd use the general form \[(x - h)^2 = 4a(y - k)\] (h, k) is the vertex and a is the focal length, a , ... dirextrix is y = k - a anf focus is (h, k + a) to find a,\[6y + 11 = 3(x - 1)^2\] so \[2y + \frac{11}{3} = (x -1)^2\] or \[2(y + \frac{11}{6}) = (x - 1)^2\] lastly you need the focal length a you know 4a = 2 so a = 1/2 now you can find the focus and directrix

OpenStudy (vegan555):

thank you so much!!